An object moving north with an initial velocity of 14 m/s accelerates 5m/s for 20 seconds what is the final velocity of the object


1
Expert's answer
2020-10-12T03:41:27-0400

The velocity of an object with constant acceleration is the following:


"v = v_0 + at"

where "v_0 = 14m\/s" is the initial veloctiy, "a = 5m\/s^2" is its acceleration, "t = 20s" is the time of acceleration. Thus, have:


"v = v_0 + at = 14m/s + 5m/s^2\cdot 20s = 114 m/s"

Answer. 114 m/s.